OpenStudy (anonymous):

H11P1: LC Tank Parallel resonant "tank" circuits are common in radio equipment. But unfortunately there is always resistance that prevents them from being perfect: in every real inductor, the wire that makes up the inductance has some resistance, and there may be leakage in the capacitor that can be modeled as a resistance. Also, the Norton resistance of the system connected to the tank circut looks like a leakage through the capacitor. So a realistic model for a tank circuit is the following: In the space provided below write an algebraic expression in terms of the device parameters for t

OpenStudy (anonymous):

first z=((RcSL+RcRl)/RcCL)/(s^2+s((L+RcRlC)/RcCL))+((Rc+Rl)/(RcCL)) so, DeltaW=((L+RcRlC)/RcCL)) simp...... DeltaW = (1/RcC)+(Rl/L) and wo^2=((Rc+Rl)/(RlCL)) remember that: wo=2pif good luck..